# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v1

@@ 7 Propoosed Algorithm and Final Results
Input: Policy $\pi$, Initial values $V_{0}$, $g_{0}=0$, Stepsizes $\eta_{t}$, $\beta_{t}$, Truncation level $N_{\max}$, $t=0,1,\ldots,T-1$, Anchor state $s_{0}\in\mathcal{S}$
@@
3:         if Contamination uncertainty set then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to (22)
4:         else if TV distance or Wasserstein distance uncertainty set then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to Algorithm 1
@@
10:         if Contamination uncertainty set then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to (22)
11:         else if TV distance or Wasserstein distance uncertainty set then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to Algorithm 1
@@
To derive the sample complexity of robust policy evaluation, we utilize the span semi-norm contraction property of the bellman operator in Theorem 2, and fit Algorithm 2 into the general biased stochastic approximation result in Theorem 3 while incorporating the bias analysis characterized in Section 6. Since each phase of Algorithm 2 contains a loop of length $T$ with all the states and actions updated together, the total samples needed for the entire algorithm in expectation is $2SAT\mathbb{E}[N_{\max}]$, where $\mathbb{E}[N_{\max}]$ is one for contamination uncertainty sets and is $\operatorname*{\mathcal{O}}(N_{\max})$ from Theorem 4 for TV and Wasserstein distance uncertainty sets.
@@
If $V_{t}$ is generated by Algorithm 2 and satisfying Assumption 1, then if the stepsize $\eta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$, we require a sample complexity of $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for contamination uncertainty set and a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for TV and Wasserstein distance uncertainty set to ensure an $\epsilon$ convergence of $V_{T}$.
@@
If $g_{t}$ is generated by Algorithm 2 and satisfying Assumption 1, then if the stepsize $\beta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$, we require a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for contamination uncertainty set and a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for TV and Wasserstein distance uncertainty set to ensure an $\epsilon$ convergence of $g_{T}$.
The formal version of Theorem 7 and Theorem 8 along with the proofs are in Appendix D. Theorem 7 and Theorem 8 provide the order-optimal sample complexity of $\tilde{\operatorname*{\mathcal{O}}}(\epsilon^{-2})$ for Algorithm 2 to achieve an $\epsilon$-accurate estimate of $V_{T}$ and $g_{T}$. The proof of Theorem 8 extends the analysis of Theorem 7 to robust average reward estimation. The key difficulty lies in controlling the propagation of error from value function estimates to reward estimation. By again leveraging the contraction property and appropriately tuning step sizes, we establish an $\tilde{\operatorname*{\mathcal{O}}}(\epsilon^{-2})$ complexity bound for robust average reward estimation.
@@ 8 Conclusion
A crucial step in our analysis is proving that the robust Bellman operator is contractive under the span semi-norm, ensuring the validity of stochastic approximation updates. We further develop a truncated multi-level Monte Carlo estimator that efficiently computes worst-case value functions under total variation and Wasserstein uncertainty, while keeping bias and variance controlled. Our results confirm that robust policy evaluation can be achieved with near-optimal efficiency, comparable to standard stochastic approximation methods.
@@ 9 Acknowledgement
Appendix A Span Semi-Norm Contraction Property of Bellman Operator
A.1 Proof of Lemma 1
Let $V_{1},V_{2}:\mathcal{S}\to\mathbb{R}$ and define $\Delta=V_{1}-V_{2}$. Denote $\mathsf{P}^{\pi}$ as the transition matrix of the Markov chain under policy $\pi$. By the definition of the Bellman operator under the transition $\mathsf{P}^{\pi}$, we have
@@ A.1 Proof of Lemma 1
Hence
$\bigl{\|}\mathbf{T}_{g}^{\pi}(V_{1})-\mathbf{T}_{g}^{\pi}(V_{2})\bigr{\|}_{\mathrm{sp}}=\|\mathsf{P}^{\pi}\,\Delta\|_{\mathrm{sp}}.$ (33)
Note that irreducibility and aperiodicity of the Markov chain on a finite state space implies (Norris, 1998) the existence of an integer $m\geq 1$ and $\gamma>0$ such that
$(\mathsf{P}^{\pi})^{m}(s_{1},s_{2})\geq\gamma\quad\forall\,s_{1},s_{2}\in\mathcal{S}.$ (34)
Thus each row of $(\mathsf{P}^{\pi})^{m}$ is a probability distribution with at least positive mass $\gamma$ over all states. We now introduce the following lemma to characterize the contraction property of $(\mathsf{P}^{\pi})^{m}$,
@@
Let $M$ be an $S\times S$ stochastic matrix with the property
$M(i,t)\;\geq\;\gamma\quad\text{for all }i,t\in\{1,\dots,S\}\quad\text{and someconstant }0<\gamma\leq 1/S.$
Then for any function $x:\{1,\dots,S\}\to\mathbb{R}$, the span of $Mx$ is strictly reduced:
$\|Mx\|_{\mathrm{sp}}\leq(1-2\,\gamma)\,\|x\|_{\mathrm{sp}},$ (35)
where
$\|x\|_{\mathrm{sp}}=\max_{1\leq s\leq S}\,x(s)-\min_{1\leq s\leq S}\,x(s).$ (36)
@@
We first define
$x_{\max}=\max_{s}x(s),\quad x_{\min}=\min_{s}x(s),\quad\|x\|_{\mathrm{sp}}=x_{\max}-x_{\min}.$
If $\|x\|_{\mathrm{sp}}=0$, the statement is trivial since both sides are zero. Otherwise, consider the normalized function
$y(s)\coloneqq\frac{x(s)-x_{\min}}{\,x_{\max}-x_{\min}\,},$
so that $0\leq y(s)\leq 1$ for all $s$ and $\|y\|_{\mathrm{sp}}=1$. Then denote $e$ as the all-ones vectors and we have
$\|Mx\|_{\mathrm{sp}}$ $=\|x_{\min}\cdot Me+(x_{\max}-x_{\min})M\bigl{[}\tfrac{x-x_{\min}}{x_{\max}-x_{\min}}\bigr{]}\|_{\mathrm{sp}}$ (37)
$=(x_{\max}-x_{\min})\bigl{\|}My\bigr{\|}_{\mathrm{sp}}$
$=\|x\|_{\mathrm{sp}}\|My\|_{\mathrm{sp}}.$
It remains to prove $\|My\|_{\mathrm{sp}}\leq(1-2\gamma)$ for $y$ taking the values in $[0,1]$.
Assume $y:\{1,\dots,S\}\to[0,1]$ with $\max_{s}y(s)-\min_{s}y(s)=1$, so in fact $\min_{s}y(s)=0$ and $\max_{s}y(s)=1$. For $1\leq i\leq S$, define
$(My)(i)=\sum_{t=1}^{S}M(i,t)y(t).$ (38)
Let $i^{*}\coloneqq\arg\max_{i}(My)(i)$ and $j^{*}\coloneqq\arg\min_{j}(My)(j)$. Then
$\|My\|_{\mathrm{sp}}=(My)(i^{*})-(My)(j^{*}).$ (39)
To bound this difference, construct a pair of random variables $(T,T^{\prime})$ as follows:
$T\sim M(i^{*},\cdot),\quad T^{\prime}\sim M(j^{*},\cdot),$ (40)
However, we can couple them in such a way that
$\mathbb{P}[T=T^{\prime}]\geq 2\gamma.$
Such a coupling is possible because each row $M(i^{*},t)$ and $M(j^{*},t)$ has at least $\gamma$ mass on every $t$. Under this coupling:
$(My)(i^{*})=\mathbb{E}\bigl{[}y(T)\bigr{]},\quad(My)(j^{*})=\mathbb{E}\bigl{[}y(T^{\prime})\bigr{]}.$
Hence
$(My)(i^{*})-(My)(j^{*})=\mathbb{E}\bigl{[}y(T)-y(T^{\prime})\bigr{]}.$ (41)
But $y(t)\in[0,1]$. Whenever $T=T^{\prime}$, the difference $y(T)-y(T^{\prime})$ is zero. Since $\text{Pr}\left[[\right]T=T^{\prime}]\geq 2\gamma$, the event $T\neq T^{\prime}$ has probability at most $(1-2\gamma)$. On $T\neq T^{\prime}$, the difference $|y(T)-y(T^{\prime})|$ is at most $1$ (because $0\leq y(\cdot)\leq 1$). Therefore
$\bigl{|}(My)(i^{*})-(My)(j^{*})\bigr{|}=\bigl{|}\mathbb{E}\bigl{[}y(T)-y(T^{\prime})\bigr{]}\bigr{|}\leq(1-2\gamma)\cdot 1=1-2\gamma.$ (42)
Thus $\|My\|_{\mathrm{sp}}\leq 1-2\gamma$ whenever $\|y\|_{\mathrm{sp}}=1$ and $y\in[0,1]$. ∎
Directly applying $(\mathsf{P}^{\pi})^{m}$ to Lemma 3 would obtain
$\|(P^{\pi})^{m}\Delta\|_{\mathrm{sp}}\leq(1-2\gamma)\|\Delta\|_{\mathrm{sp}},$ (43)
Where $\gamma$ is defined in (34). Thus there also exist $\beta\in(0,1)$ where $\beta=(1-2\gamma)^{\frac{1}{m}}$ such that $\|(P^{\pi})\Delta\|_{\mathrm{sp}}\leq\beta\|\Delta\|_{\mathrm{sp}}$. Thus, this proves the lemma.
A.2 Proof of Theorem 2
We note that for any $g$, for any $V_{1}$ and $V_{2}$, for all $s\in\mathcal{S}$, we have:

### v2

@@ 5 Robust Average Reward TD Learning
Input: Policy $\pi$, Initial values $V_{0}$, $g_{0}=0$, Stepsizes $\eta_{t}$, $\beta_{t}$, Max level $N_{\max}$, Anchor state $s_{0}\in\mathcal{S}$
@@
3:         if Contamination then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to (18)
4:         else if TV or Wasserstein then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to Algorithm 1
@@
10:         if Contamination then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to (18)
11:         else if TV or Wasserstein then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to Algorithm 1
@@
To derive the sample complexity of robust policy evaluation, we utilize the semi-norm contraction property of the Bellman operator in Theorem 2, and fit Algorithm 2 into the general biased stochastic approximation result in Theorem 8 while incorporating the bias analysis characterized in Section 4. Since each phase of Algorithm 2 contains a loop of length $T$ with all the states and actions updated together, the total samples needed for the entire algorithm in expectation is $2SAT\mathbb{E}[N_{\max}]$, where $\mathbb{E}[N_{\max}]$ is one for contamination uncertainty sets and is $\operatorname*{\mathcal{O}}(N_{\max})$ from Theorem 3 for TV and Wasserstein distance uncertainty sets.
@@
If $V_{t}$ is generated by Algorithm 2 and satisfying Assumption 1, then if the stepsize $\eta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$, we require a sample complexity of $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for contamination uncertainty set and a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for TV and Wasserstein distance uncertainty set to ensure an $\epsilon$ convergence of $V_{T}$. Moreover, these results are order-optimal in terms of $\epsilon$.
@@
If $g_{t}$ is generated by Algorithm 2 and satisfying Assumption 1, then if the stepsize $\beta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$, we require a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for all contamination, TV, and Wasserstein distance uncertainty set to ensure an $\epsilon$ convergence of $g_{T}$.
The formal version of Theorems 6 and 7 along with the proofs are in Appendix F. Theorem 6 provides the order-optimal sample complexity of $\tilde{\operatorname*{\mathcal{O}}}(\epsilon^{-2})$ for Algorithm 2 to achieve an $\epsilon$-accurate estimate of $V_{T}$. The proof of Theorem 7 extends the analysis of Theorem 6 to robust average reward estimation. The key difficulty lies in controlling the propagation of error from value function estimates to reward estimation. By again leveraging the contraction property and appropriately tuning step sizes, we establish an $\tilde{\operatorname*{\mathcal{O}}}(\epsilon^{-2})$ complexity bound for robust average reward estimation.
@@ 6 Conclusion
A crucial step in our analysis is proving that the robust Bellman operator is contractive under our constructed semi-norm $\|\cdot\|_{\mathcal{P}}$, ensuring the validity of stochastic approximation updates. We further develop a truncated multi-level Monte Carlo estimator that efficiently computes worst-case value functions under total variation and Wasserstein uncertainty, while keeping bias and variance controlled. One limitation of this work is that the results require ergodicity to hold in the setting, as stated in Assumption 1. Additionally, scaling the algorithm and results in the paper via function approximations remains an important open problem.
@@ 7 Acknowledgement
Appendix A Related Work
The theoretical guarantees of robust average reward have been studied by the following works. Wang et al. (2023a) takes a model-based perspective, approximating robust average-reward MDPs with discounted MDPs and proving uniform convergence of the robust discounted value function as the discount factor approaches one, employing dynamic programming and Blackwell optimality arguments to characterize optimal policies. Wang et al. (2023b) propose a model-free approach by developing robust relative value iteration (RVI) TD and Q-learning algorithms, proving their almost sure convergence using stochastic approximation, martingale theory, and multi-level Monte Carlo estimators to handle non-linearity in the robust Bellman operator. While these studies provide fundamental insights into robust average-reward RL, they do not establish explicit convergence rate guarantees due to the lack of contraction properties in the robust Bellman operator. In addition, Sun et al. (2024) studies the policy optimization of average reward robust MDPs with assuming direct queries of the value functions.
Policy evaluation in robust discounted-reward reinforcement learning with finite sample guarantees has been extensively studied, with the key recent works Wang and Zou (2022); Zhou et al. (2024); Li et al. (2022); Kumar et al. (2023); Kuang et al. (2022) focusing on solving the robust Bellman equation by finding its fixed-point solution. This approach is made feasible by the contraction property of the robust Bellman operator under the sup-norm, which arises due to the presence of a discount factor $\gamma<1$. However, this fundamental approach does not directly extend to the robust average-reward setting, where the absence of a discount factor removes the contraction property under any norm. As a result, existing robust discounted methods cannot be applied in the robust average reward RL setting.
Appendix B Relaxing Assumption 1
In this section, we provide further justifications on Remark 1 and characterize the conditions when Assumption 1 could be relaxed into only assuming the nominal transition being irreducible and aperiodic for contamination, TV distance and Wasserstein distance uncertainty sets.
Lemma 3.
Let $\pi$ be a stationary policy on a finite state space $S$, and let $P_{0}$ be the nominal kernel under $\pi$ whose induced chain is irreducible and aperiodic. Define
$m_{\min}=\min_{\begin{subarray}{c}s,a,s^{\prime}\\ P_{0}(s^{\prime}|s,a)>0\end{subarray}}P_{0}(s^{\prime}|s,a)>0,\qquad d_{\min}=\min_{x\neq y}d(x,y)>0$
where $d$ is the metric for the Wasserstein uncertainty set. Then there exists $\bar{\delta}>0$ such that for every $\delta\in[0,\bar{\delta})$, any $P$ in each of the following families remains irreducible and aperiodic under $\pi$:
$\mathcal{P}_{\rm cont}(\delta)=\Bigl{\{}P:P=(1-\delta)P_{0}+\delta\,Q,\;Q\in\Delta(S)\Bigr{\}},$
$\mathcal{P}_{\rm TV}(\delta)=\Bigl{\{}Q\in\Delta(S):\tfrac{1}{2}\|Q-P_{0}\|_{1}\leq\delta\Bigr{\}},$
$\mathcal{P}_{W}(\delta)=\Bigl{\{}Q\in\Delta(S):W_{1}(Q,P_{0})\leq\delta\Bigr{\}}.$
Moreover one may take
$\bar{\delta}=\min\Bigl{\{}1,\;\tfrac{1}{2}\,m_{\min},\;m_{\min}\,d_{\min}\Bigr{\}}.$
Proof.
We show that for each uncertainty type, whenever $\delta<\bar{\delta}$ every originally positive transition under $P_{0}$ remains strictly positive in $P$. Preservation of the support–graph and any self–loops then guarantees irreducibility and period one.
1. Contamination. If
$P(s^{\prime}|s,a)=(1-\delta)\,P_{0}(s^{\prime}|s,a)+\delta\,Q(s^{\prime}|s,a),$
then
$P(s^{\prime}\!\mid s,a)\geq(1-\delta)P_{0}(s^{\prime}\!\mid s,a).$
Since $P_{0}(s^{\prime}\!\mid s,a)>0$ and $\delta<1$, each positive edge of $P_{0}$ survives, as do any self–loops.
2. Total‐variation. If $\tfrac{1}{2}\|Q-P_{0}\|_{1}\leq\delta$, then
$|Q(s^{\prime})-P_{0}(s^{\prime})|\leq 2\delta,\quad\Longrightarrow\quad Q(s^{\prime})\geq P_{0}(s^{\prime})-2\delta.$
Define $m_{\min}=\min\{P_{0}(s^{\prime}\!\mid s,a)>0\}$. For $\delta<\tfrac{1}{2}\,m_{\min}$, we have $Q(s^{\prime}\!\mid s,a)\geq m_{\min}-2\delta>0$. Thus all original transitions remain.
3. Wasserstein. If $W_{1}(Q,P_{0})\leq\delta$, moving mass $P_{0}(s^{\prime})$ from $s^{\prime}$ to any other state costs at least
$P_{0}(s^{\prime})\,d_{\min}.$
Hence $Q(s^{\prime})=0$ would force $W_{1}(Q,P_{0})\geq P_{0}(s^{\prime})\,d_{\min}$. Taking $\delta<m_{\min}\,d_{\min}$ prevents zeroing any originally positive entry.
In all three cases, the perturbed kernel $P$ retains exactly the positive‐support and self–loops of $P_{0}$. Therefore irreducibility and aperiodicity hold uniformly over $\mathcal{P}_{\rm cont}(\delta)$, $\mathcal{P}_{\rm TV}(\delta)$, and $\mathcal{P}_{W}(\delta)$ whenever $\delta<\bar{\delta}$. ∎
Appendix C Semi-Norm Contraction Property of the Bellman Operator
C.1 Proof of Lemma 1
For any $V_{1},V_{2}\in\mathbb{R}^{S}$ and define $\Delta=V_{1}-V_{2}$. Denote $\mathsf{P}^{\pi}$ as the transition matrix under policy $\pi$ and the unique stationary distribution $d^{\pi}$, and denote $E$ as the matrix with all rows being identical to $d^{\pi}$. We further define $Q^{\pi}=\mathsf{P}^{\pi}-E$. Thus, we would have,
@@ C.1 Proof of Lemma 1
which implies
$\mathbf{T}_{g}^{\pi}(V_{1})-\mathbf{T}_{g}^{\pi}(V_{2})=\mathsf{P}^{\pi}\Delta=Q^{\pi}\Delta+E\Delta$ (27)
We now discuss the detailed construction of the semi-norm $\|\cdot\|_{\mathsf{P}}$. Since $\mathsf{P}^{\pi}$ is ergodic, according to the Perron–Frobenius theorem, $\mathsf{P}^{\pi}$ has an eigenvalue $\lambda_{1}=1$ of algebraic multiplicity exactly one, with corresponding right eigenvector $\mathbf{e}$. Moreover, all other eigenvalues $\lambda_{2}\geq\ldots\geq\lambda_{S}$ of $\mathsf{P}^{\pi}$ satisfies $|\lambda_{i}|<1$ for all $i\in\{2,\ldots,S\}$.
@@
All eigenvalues of $Q^{\pi}$ lies strictly inside the unit circle.
@@
Since $E=\mathbf{e}d^{\pi\top}$, $E$ is a rank‑one projector onto the span of $\mathbf{e}$. Hence the spectrum of $E$ is $\{1,0,\ldots,0\}$. In addition, we can show $\mathsf{P}^{\pi}$ and $E$ commute by
$\mathsf{P}^{\pi}E=\mathsf{P}^{\pi}(\mathbf{e}d^{\pi})=(\mathbf{e}d^{\pi})=E$ (28)
$E\mathsf{P}^{\pi}=\mathbf{e}(d^{\pi}\mathsf{P}^{\pi})=\mathbf{e}d^{\pi}=E$ (29)
Thus, by the Schur theorem, $\mathsf{P}$ and $E$ are simultaneously upper triangularizable. In a common triangular basis, the diagonals of $\mathsf{P}$ and $E$ list their eigenvalues in descending orders, which are $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{S}\}$ and $\{1,0,\ldots,0\}$ respectively. Thus, in that same basis, $Q^{\pi}=\mathsf{P}^{\pi}-E$ is also triangular, with diagonal entries being $\{\lambda_{1}-1,\lambda_{2}-0,\ldots,\lambda_{S}-0\}$. Since $\lambda_{1}=1$, we have the spectrum of $Q^{\pi}$ is exactly $\{\lambda_{2},\ldots,\lambda_{S},0\}$. Since we already have $|\lambda_{i}|<1$ for all $i\in\{2,\ldots,S\}$, we conclude the proof. ∎
Define $\rho(\cdot)$ to be the spectral radius of a matrix, then Lemma 4 implies that $\rho(Q^{\pi})<1$. Hence by equivalence of norms in $\mathbb{R}^{|\mathcal{S}|}$ it is possible to construct a vector norm $\|\cdot\|_{Q}$ so that the induced operator norm of $Q^{\pi}$ is less than $1$, specifically
$10\leq\rho(Q^{\pi})\leq\bigl{\|}Q^{\pi}\bigr{\|}_{Q}\leq\alpha<$ 1 . (30)
A concrete construction example is to leverage the discrete-Lyapunov equation Horn and Johnson (2012) of solving $M$ on the space of symmetric matrices for any $\rho(Q^{\pi})<\alpha<1$ as follows:
$Q^{\pi\top}MQ^{\pi}-\alpha^{2}M=-I$ (31)
We can express $M$ in the form of Neumann‐series as
$M$ $=\alpha^{-2}I+\alpha^{-4}(Q^{\pi})^{\top}Q^{\pi}+\alpha^{-6}\left((Q^{\pi})^{\top}\right)^{2}(Q^{\pi})^{2}+\dots$
$=\sum_{k=0}^{\infty}\alpha^{-2(k+1)}\left((Q^{\pi})^{\top}\right)^{k}(Q^{\pi})^{k}.$ (32)
Since $\rho((Q^{\pi})^{\top}Q^{\pi})=\rho(Q^{\pi})^{2}<\alpha^{2}<1$, the partial sum expression in (C.1) converges absolutely as a geometric‐type series. Also, since each term in (C.1) is positive semi‑definite, and the first term $\alpha^{-2}I$ being positive definite, we can conclude that $M$ being the summation is well-defined and is positive definite. Thus, using the $M$ defined in (31), we can define our desired $\|\cdot\|_{Q}$ as
$\|x\|_{Q}\coloneqq\sqrt{x^{\top}Mx}$ (33)
which implies
$\|Q^{\pi}\|_{Q}=\sup_{x\neq\bf{0}}\frac{\|Q^{\pi}x\|_{Q}}{\|x\|_{Q}}=\sup_{x\neq\bf{0}}\frac{\sqrt{(Q^{\pi}x)^{\top}MQ^{\pi}x}}{\sqrt{x^{\top}Mx}}\overset{(a)}{\leq}\alpha<1$ (34)
Where $(a)$ is because for any $x\neq\bf{0}$, from (31) we have
$(Qx)^{\pi\top}MQ^{\pi}x-\alpha^{2}x^{\top}Mx=-x^{\top}x\quad\Rightarrow\quad(Qx)^{\pi\top}MQ^{\pi}x=\alpha^{2}x^{\top}Mx-\|x\|_{2}^{2}$ (35)
Since $\|x\|_{2}^{2}$ is always non-negative dividing both sides of the second equation of (35) by $x^{\top}Mx$ and further taking the square root on both sides yields the inequality of $(a)$.
Based on the above construction of $\|\cdot\|_{Q}$, define the operator $\|\cdot\|_{\mathsf{P}}$ as
$\|x\|_{\mathsf{P}}\coloneqq\bigl{\|}Q^{\pi}x\bigr{\|}_{Q}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{Q}$ (36)
where $0<\epsilon<1-\alpha$.
Lemma 5.
The operator $\|\cdot\|_{\mathsf{P}}$ is a valid semi-norm with kernel being exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. Furthermore, for all $x\in\mathbb{R}^{S}$, we have $\bigl{\|}\mathsf{P}^{\pi}x\bigr{\|}_{\mathsf{P}}\leq(\alpha+\epsilon)\bigl{\|}x\bigr{\|}_{\mathsf{P}}$.
Proof.
Regarding positive homogeneity and nonnegativity, for any scalar $\lambda$ and $x\in\mathbb{R}^{S}$,
$\|\lambda x\|_{\mathsf{P}}=\bigl{\|}Q^{\pi}(\lambda x)\bigr{\|}_{Q}+\epsilon\inf_{c}\bigl{\|}\lambda x-c\mathbf{e}\bigr{\|}_{Q}=|\lambda|\bigl{\|}Q^{\pi}x\bigr{\|}_{Q}+\epsilon\,|\lambda|\inf_{c}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{Q}=|\lambda|\,\|x\|_{\mathsf{P}},$
and clearly $\|x\|_{\mathsf{P}}\geq 0$, with equality only when both $\|Q^{\pi}x\|_{Q}=0$ and $\inf_{c}\|x-c\,e\|_{Q}=0$. Regarding triangle inequality, for any $x,y\in\mathbb{R}^{S}$,
$\|x+y\|_{\mathsf{P}}$ $=\bigl{\|}Q^{\pi}(x+y)\bigr{\|}_{Q}+\epsilon\inf_{c}\bigl{\|}x+y-c\mathbf{e}\bigr{\|}_{Q}$
$\leq\bigl{\|}Q^{\pi}x\bigr{\|}_{Q}+\bigl{\|}Q^{\pi}y\bigr{\|}_{Q}+\epsilon\inf_{a,b}\bigl{\|}x-a\mathbf{e}+y-b\mathbf{e}\bigr{\|}_{Q}$
$\leq\bigl{\|}Q^{\pi}x\bigr{\|}_{Q}+\bigl{\|}Q^{\pi}y\bigr{\|}_{Q}+\epsilon\inf_{a}\bigl{\|}x-a\mathbf{e}\bigr{\|}_{Q}+\epsilon\inf_{b}\bigl{\|}y-b\mathbf{e}\bigr{\|}_{Q}$
$=\|x\|_{\mathsf{P}}+\|y\|_{\mathsf{P}}.$
Regarding the kernel, if $x=k\mathbf{e}$ for some $k\in\mathbb{R}$, then we have
$\|x\|_{\mathsf{P}}$ $=\|kQ^{\pi}\mathbf{e}\|_{Q}+\epsilon\inf_{c}\|k\mathbf{e}-c\mathbf{e}\|_{Q}$
$=\|k(\mathsf{P}^{\pi}-E)\mathbf{e}\|_{Q}+\epsilon\|k\mathbf{e}-k\mathbf{e}\|_{Q}$
$=\|k\mathbf{e}-k\mathbf{e}\|_{Q}+\epsilon\|0\|_{Q}=0$ (37)
On the other hand, if $x\notin\{c\mathbf{e}:c\in\mathbb{R}\}$, we know that
$\|x\|_{\mathsf{P}}\geq\epsilon\inf_{c}\|x-c\mathbf{e}\|_{Q}>0$ (38)
Thus, the kernel of $\|\cdot\|_{\mathsf{P}}$ is exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. We now show that, for any $x\in\mathbb{R}^{S}$,
$\bigl{\|}Q^{\pi}x\bigr{\|}_{\mathsf{P}}$ $=\bigl{\|}Q^{\pi}(Q^{\pi}x)\bigr{\|}_{Q}+\epsilon\inf_{c}\bigl{\|}Q^{\pi}x-c\mathbf{e}\bigr{\|}_{Q}$
$\leq\alpha\|Q^{\pi}x\|_{Q}+\epsilon\|Q^{\pi}x\|_{Q}$
$=(\alpha+\epsilon)\|Q^{\pi}x\|_{Q}$
$\leq(\alpha+\epsilon)\|x\|_{\mathsf{P}}.$ (39)
Let $\beta=\alpha+\epsilon$, by (30) and (36), we have $\alpha\in(0,1)$ and $\epsilon\in(0,1-\alpha)$. Thus, $\beta\in(0,1)$ and combining $\beta$ with the semi-norm $\|\cdot\|_{\mathsf{P}}$ confirms Lemma 1.
C.2 Proof of Theorem 2
We override the terms $\alpha,\lambda$ and $\epsilon$ from the previous section. For any $V_{1},V_{2}$ and $s\in\mathcal{S}$,